The pressure of a given mass of gas is tripled and the temperature in
kelvin scale is doubled. Find the percentage change in volume.
1
Expert's answer
2020-12-30T04:05:53-0500

"\frac {P1V1}{T1}=\frac {P2V2}{T2}"


"\frac {P1V1}{T1}=\frac {3P1V1}{3T1}=\frac {1.5P2V2}{T2}"

Therefore, V2 went up by 50%.